OpenStudy (anonymous):

how would i graph .7x+75 ?

OpenStudy (anonymous):

this is y = .7x + 75. Do you have a graphing calculator? You can create a table of values for x and y values. Plug in each x value into that equation to get its respective y coordinate. This graph will be a straight line that will intercept the y -axis at 75 (which is when x = 0)....positive slope.

OpenStudy (anonymous):

well since its a linear equation, we conclude that it is a straight line. Now to plot a straight line we either need two points on line or a point and a slope. put x=0 in equation u get y=75 (a point on line). now for slope. we know this y=mx+a where m is the slope. so here slope becomes .7, u have a point and slope.. thats all u need

OpenStudy (anonymous):

A plot is attached

OpenStudy (anonymous):

i get how i let x=0 so y becomes 75. But is it also correct if i let y=0 and x would be .7x=-75? just in case i want another point on the line besides (0,75) Thanks everyone

OpenStudy (anonymous):

y=(7/10) x + 75, When x=0, y=+75 Solving for x, x=(10/7) (y - 75) When y=0, x= -750/7 = -107.1 Refer to the second version of the plot. It shows the x and y axis and the associated intercepts. Points (0,75) and (-107.1,0) determine your line and you can see these points on the attached plot.

OpenStudy (anonymous):

@trigger yaa thats correct.. that would another point on line (called x-intercept).

OpenStudy (anonymous):

Thanks everyone, my calculus test is tomrrow and its 3:44 am here. Kind of screwed. Summer school sucks

OpenStudy (anonymous):

Good Luck on your test and thank you for the medal